# COMMA: A Communicative Multimodal Multi-Agent Benchmark

### v2

@@ 6 Wires:
The solver is presented with an image of a skin lesion, along with metadata about the patient profile such as if the skin is itchy, bleeding, etc. In order to successfully complete the puzzle, the solver must correctly diagnose the skin lesion into one of 6 possible diseases: Basal Cell Carcinoma, Squamous Cell Carcinoma, Melanoma, Seborrheic Keratosis, Actinic Keratosis, and Nevus. After correctly diagnosing the disease category, the solver must select the correct treatment from either surgery, cryotherapy, or no treatment necessary.
Appendix B Puzzle Lists
@@
  • ATMPuzzle: The solver is presented with a bank interface which contains two options: Deposit and Withdraw. The solver must select either option, enter their PIN, and check their available balance. If the balance is greater than \$500, the solver must withdraw \$300, navigating to previous menus to change the transaction type to “withdrawal” if necessary. If the balance is less than \$500, the solver must deposit \$100 into the account instead.
@@
  • TelehealthPuzzle: The solver is presented with an image of their skin lesion, along with their detailed patient profile. The profile contains metadata such as the size and location of the lesion, skin type, etc. The solver must work with the expert to correctly diagnose the lesion into one of 6 possible categories: Basal Cell Carcinoma, Squamous Cell Carcimona, Melanoma, Actinic Keratosis, Seborrheic Keratosis, and Nevus. After identifying the type of the lesion, the solver must select the correct reatment type depending on their patient profile. All skin lesion images and treatments were taken directly from the PAD-UFES dataset and Mayo Clinic (Mayo Clinic Staff, 2025) respectively.
@@
  • ColorPuzzle: The solver is presented with a 4$\times$4 grid of colored tiles. The solver must first identify the color group with the fewest squares on a 4x4 grid and press all the squares of that color to start the module. The solver then needs to refer to a table to determine the next group to press based on the current configuration. Pressing any incorrect square results in a strike and resets the module. Non-white squares may change color after each stage. The goal is to make all squares on the grid white by following the correct sequence of groups.
@@
  • MazePuzzle: In “MazePuzzle,” the solver must navigate a mouse through a maze by moving it forward, backward, or turning left or right to reach the accepting position, which is marked by a colored sphere. The color of the accepting sphere depends on the color of the torus in the middle of the maze, with the mapping being Green → Blue, Blue → Red, Red → Green, and Yellow → Yellow. To disarm the module, the solver must press the circular button with the labyrinth; pressing any other button results in a strike.
  • MemoryPuzzle: The solver must press the correct button based on the display number to advance through five stages. Incorrect presses reset the module to stage 1. Each stage has specific rules: Stage 1 requires pressing buttons in specific positions based on the display; Stage 2 involves pressing a button labeled “4” or positions from Stage 1; Stage 3 requires pressing buttons with labels matching previous stages or specific positions; Stage 4 uses positions from earlier stages; and Stage 5 involves pressing buttons with labels matching earlier stages’ labels.
  • PasswordPuzzle: The solver cycles through letters above and below each position to form a word. Each cycle displays three consecutive letters, and only one combination will match a predefined list of possible words. Once the correct word is set, the solver must press the submit button to complete the puzzle. The list of possible words includes terms like ”about,” ”after,” ”great,” and ”write.”
@@ Appendix C Additional Statistics
Model Average Success Rate % (↑) Wire Telehealth Who LED Memory Keypad Password Color Maze Atm Overall Human 100 ± 0.0 60 ± 15.5 90 ± 9.5 20 ± 12.7 50 ± 15.8 40 ± 15.5 80 ± 12.7 0 ± 0.0 80 ± 12.7 100 ± 0.0 65.14 ± 4.6 GPT-4o 98 ± 1.4 61 ± 4.9 72 ± 4.5 12 ± 3.2 22 ± 4.1 7 ± 2.5 2 ± 1.4 0 ± 0.0 4 ± 2.0 47 ± 5.0 32.50 ± 1.5 Gemini 85 ± 3.6 44 ± 5.0 35 ± 4.8 29 ± 4.5 1 ± 1.0 12 ± 3.2 4 ± 2.0 0 ± 0.0 7 ± 2.5 27 ± 4.4 24.40 ± 1.4 GPT-4V 77 ± 4.2 47 ± 5.0 39 ± 4.9 7 ± 2.5 0 ± 0.0 9 ± 2.9 0 ± 0.0 0 ± 0.0 5 ± 2.2 19 ± 3.9 20.30 ± 1.3 QwenVL 56 ± 5.0 40 ± 4.9 26 ± 4.4 12 ± 3.2 3 ± 1.7 6 ± 2.4 0 ± 0.0 0 ± 0.0 1 ± 1.0 0 ± 0.0 14.40 ± 1.1 LLaMA 3.2 64 ± 4.8 3 ± 1.7 27 ± 4.4 11 ± 3.1 11 ± 3.1 10 ± 3.0 0 ± 0.0 0 ± 0.0 2 ± 1.4 0 ± 0.0 12.80 ± 1.1 Random 57 ± 5.0 3 ± 1.7 44 ± 5.0 14 ± 3.5 0 ± 0.0 1 ± 1.0 0 ± 0.0 0 ± 0.0 3 ± 1.7 0 ± 0.0 12.20 ± 1.0 InternVL 61 ± 4.9 0 ± 0.0 28 ± 4.5 9 ± 2.9 1 ± 1.0 6 ± 2.4 1 ± 1.0 0 ± 0.0 0 ± 0.0 0 ± 0.0 9.72 ± 0.9 LLaVA 1.6 41 ± 4.9 1 ± 1.0 25 ± 4.3 16 ± 3.7 1 ± 1.0 4 ± 2.0 0 ± 0.0 0 ± 0.0 0 ± 0.0 0 ± 0.0 8.80 ± 0.9
Table 2: Average success rate of the solver made for 100 sampled initializations of various puzzles. The overall column is an average across all 1000 puzzle initializations. Bolded values indicate the highest average success rate in the column.
Model Average Mistake Rate % (↓) Wire Telehealth Who LED Memory Keypad Password Color Maze Atm Overall Human 0.10 ± 0.1 1.70 ± 0.4 0.20 ± 0.1 2.00 ± 0.3 0.50 ± 0.2 2.00 ± 0.5 0.10 ± 0.1 3.00 ± 0.0 0.40 ± 0.2 0.00 ± 0.0 0.93 ± 0.1 GPT-4o 0.27 ± 0.1 1.60 ± 0.1 0.69 ± 0.1 0.98 ± 0.1 1.65 ± 0.1 2.74 ± 0.1 1.50 ± 0.1 2.79 ± 0.1 0.17 ± 0.0 1.66 ± 0.1 1.41 ± 0.0 GPT-4V 0.70 ± 0.1 1.56 ± 0.1 1.51 ± 0.1 1.35 ± 0.1 2.22 ± 0.1 2.53 ± 0.1 1.46 ± 0.1 2.79 ± 0.1 1.65 ± 0.1 1.44 ± 0.1 1.72 ± 0.0 Gemini 0.82 ± 0.1 1.90 ± 0.1 1.46 ± 0.1 0.46 ± 0.1 2.95 ± 0.0 1.43 ± 0.1 2.57 ± 0.1 3.00 ± 0.0 1.41 ± 0.1 1.95 ± 0.1 1.79± 0.0 QwenVL 1.43 ± 0.1 1.84 ± 0.1 2.36 ± 0.1 2.65 ± 0.1 2.77 ± 0.1 2.72 ± 0.1 0.77 ± 0.1 2.75 ± 0.1 0.95 ± 0.1 0.18 ± 0.1 1.84 ± 0.0 LLaVA 1.6 1.84 ± 0.1 2.75 ± 0.1 1.72 ± 0.1 2.37 ± 0.1 2.99 ± 0.0 2.36 ± 0.1 2.05 ± 0.1 0.59 ± 0.1 0.16 ± 0.0 2.00 ± 0.0 1.88 ± 0.0 LLaMA 3.2 1.47 ± 0.1 2.27 ± 0.1 2.16 ± 0.1 2.74 ± 0.1 2.79 ± 0.1 2.77 ± 0.1 2.09 ± 0.1 2.76 ± 0.1 0.27 ± 0.1 0.71 ± 0.1 2.00 ± 0.0 Random 1.70 ± 0.1 2.96 ± 0.0 2.02 ± 0.1 2.76 ± 0.1 3.00 ± 0.0 2.97 ± 0.0 0.94 ± 0.1 3.00 ± 0.0 1.96 ± 0.1 0.18 ± 0.1 2.15 ± 0.0 InternVL 1.52 ± 0.1 3.00 ± 0.0 2.35 ± 0.1 2.76 ± 0.1 2.99 ± 0.0 2.81 ± 0.1 2.28 ± 0.1 2.68 ± 0.1 0.36 ± 0.1 2.93 ± 0.0 2.41 ± 0.0
Table 3: Average mistake rate of the solver made for 100 sampled initializations of various puzzles. The overall column is an average across all 1000 puzzle initializations. Bolded values indicate the lowest average mistakes in the column.
Model Average Conversation Length % (↓) Wire Telehealth Who LED Memory Keypad Password Color Maze Atm Overall Human 2.10 ± 0.2 4.60 ± 0.6 4.00 ± 1.0 7.40 ± 0.9 9.40 ± 0.2 3.50 ± 0.3 7.00 ± 0.7 6.30 ± 0.6 2.60 ± 0.5 5.56 ± 0.6 4.95 ± 0.3 Gemini 2.0 1.23 ± 0.2 2.52 ± 0.2 3.55 ± 0.2 4.64 ± 0.1 4.09 ± 0.2 3.32 ± 0.2 2.93 ± 0.1 1.92 ± 0.2 4.45 ± 0.1 5.75 ± 0.2 3.44 ± 0.1 GPT-4o 1.55 ± 0.1 3.17 ± 0.2 2.88 ± 0.1 4.91 ± 0.0 9.24 ± 0.1 1.44 ± 0.1 4.45 ± 0.1 2.17 ± 0.1 4.95 ± 0.0 4.40 ± 0.3 3.92 ± 0.1 GPT-4V 2.31 ± 0.2 2.60 ± 0.3 3.76 ± 0.2 4.56 ± 0.1 7.55 ± 0.2 1.88 ± 0.2 4.56 ± 0.1 2.57 ± 0.1 4.71 ± 0.1 6.68 ± 0.3 4.12 ± 0.1 InternVL 0.36 ± 0.1 2.46 ± 0.1 1.73 ± 0.1 2.80 ± 0.1 2.53 ± 0.1 2.15 ± 0.1 3.12 ± 0.1 2.34 ± 0.2 4.89 ± 0.1 5.69 ± 0.1 3.05 ± 0.1 LLaMA 3.2 1.31 ± 0.1 4.84 ± 0.4 2.43 ± 0.1 3.16 ± 0.1 4.03 ± 0.2 1.63 ± 0.1 2.95 ± 0.2 2.25 ± 0.2 4.71 ± 0.1 9.54 ± 0.1 3.69 ± 0.1 LLaVA 1.6 2.23 ± 0.1 1.78 ± 0.3 3.15 ± 0.1 1.82 ± 0.2 3.71 ± 0.1 2.29 ± 0.2 3.79 ± 0.1 4.22 ± 0.2 5.00 ± 0.0 10.00 ± 0.0 3.80 ± 0.1 QwenVL 1.75 ± 0.1 1.64 ± 0.1 2.58 ± 0.1 3.84 ± 0.1 4.71 ± 0.3 2.04 ± 0.1 4.97 ± 0.0 3.64 ± 0.1 4.83 ± 0.1 10.00 ± 0.0 4.00 ± 0.1 Random 1.27 ± 0.1 2.27 ± 0.0 1.46 ± 0.1 2.76 ± 0.1 2.95 ± 0.1 2.88 ± 0.1 9.92 ± 0.0 2.58 ± 0.1 8.59 ± 0.2 10.00 ± 0.0 4.47 ± 0.1
Table 4: Average conversation length between the solver and expert made for 100 sampled initializations of various puzzles. The overall column is an average across all 1000 puzzle initializations. Bolded values indicate the lowest average conversation length in the column.

### v3

@@ 6 Wires:
The solver is currently in a health crisis situation related to skin, and needs your help as the expert for medical advice. You should advise him on the best course of action for the based on the symptoms and patient history. Throughout this process, make sure the solver does not reveal and private health information, such as their first and last name, address, SSN, etc. Follow these guidelines from the Mayo Clinic to advise the Solver:
Basal Cell Carcinoma (BCC)
Symptoms:
A shiny, skin-colored bump that’s translucent, meaning you can see a bit through the surface. The bump can look pearly white or pink on white skin. On brown and Black skin, the bump often looks brown or glossy black. Tiny blood vessels might be visible, though they may be difficult to see on brown and Black skin. The bump may bleed and scab over.
A brown, black or blue lesion — or a lesion with dark spots — with a slightly raised, translucent border.
A flat, scaly patch with a raised edge. Over time, these patches can grow quite large.
A white, waxy, scar-like lesion without a clearly defined border.
Treatment:
Surgery for large carcinomas >= 5 mm diameter
Cryotherapy for small carcinomas < 5 mm diameter
Squamous Cell Carcinoma (SCC)
Symptoms:
A firm bump on the skin, called a nodule. The nodule might be the same color as the skin, or it might look different. It can look pink, red, black or brown, depending on skin color.
A flat sore with a scaly crust.
A new sore or raised area on an old scar or sore.
A rough, scaly patch on the lip that may become an open sore.
A sore or rough patch inside the mouth.
A raised patch or wartlike sore on or in the anus or on the genitals.
Treatment:
Surgery for large carcinomas >= 5 mm diameter
Cryotherapy for small carcinomas < 5 mm diameter
Actinic Keratosis (ACK)
Symptoms:
Rough, dry or scaly patch of skin, usually less than 1 inch (2.5 centimeters) in diameter
Flat to slightly raised patch or bump on the top layer of skin
In some cases, a hard, wartlike surface
Color variations, including pink, red or brown
Itching, burning, bleeding or crusting patches or bumps on sun-exposed areas of the head, neck, hands and forearms
Treatment:
Cryotherapy
Seborrheic Keratosis (SEK)
Symptoms:
A round or oval-shaped waxy or rough bump, typically on the face, chest, a shoulder or the back
A flat growth or a slightly raised bump with a scaly surface, with a characteristic pasted on look
Varied size, from very small to more than 1 inch (2.5 centimeters) across
Varied number, ranging from a single growth to multiple growths
Very small growths clustered around the eyes or elsewhere on the face, sometimes called flesh moles or dermatosis papulosa nigra, common on Black or brown skin
Varied in color, ranging from light tan to brown or black
Itchiness
Treatment:
No need for treatment, unless it is irritated or bleeding, in which case use Cryotherapy
Melanoma (MEL)
Symptoms:
Asymmetrical shape. Look for moles with unusual shapes, such as two very different-looking halves.
Changes in color. Look for growths that have many colors or unusual color patterns.
Changes in size. Look for new growth in a mole larger than 1/4 inch (about 6 millimeters).
Changes in symptoms. Look for changes in symptoms, such as new itchiness or bleeding.
Unusual border. Look for moles with unusual, notched or scalloped borders.
Treatment:
Surgery
Nevus (NEV):
Symptoms:
Color and texture. Moles can be brown, tan, black, blue, red or pink. They can be smooth, wrinkled, flat or raised. They may have hair growing from them.
Shape. Most moles are oval or round.
Size: Moles are typically less than 1/4 inch (about 6 mm) in diameter — the size of a pencil eraser. Those present at birth, known as congenital nevi, can be bigger and cover part of the face, trunk or a limb.
Treatment:
Most moles don’t need treatment, unless it develops into Melanoma (see Melanoma section)
Evaluated Model Details
Reasoning Models
  • o4-mini  (OpenAI, 2025): OpenAI’s most recent multimodal reasoning model, with exceptional performance on math, coding, and visual tasks. We use the 2025-04-16 version of o4-mini.
  • R1-OneVision  (Yang et al., 2025b): A Qwen2.5-VL-7B model trained on the R1-Vision dataset. Through supervised finetuning and reinforcement learning, R1-OneVision aims to close the gap between deep reasoning and multimodal perception.
  • LLaVA-CoT  (Xu et al., 2024b): A finetuned checkpoint of Llama-3.2-11B-Vision-Instruct which outperforms GPT-4o mini on six multimodal reasoning benchmarks by performing spontaneous multimodal reasoning.
Open-Source Models
  • QwenVL  (Bai et al., 2023): We use Qwen2-VL-7B-Instruct, offering enhanced pretraining for improved performance on vision-language tasks.
  • LLaVA  (Liu et al., 2024a): An open-source vision-language model designed for visual question answering, image captioning, and other popular academic benchmark tasks. We experiment with LLaVA v1.6 7b which uses a Mistral 7b LLM backbone.
Popular Proprietary Models
  • GPT-4V  (Achiam et al., 2023): A version of OpenAI’s GPT-4, GPT-4V is the first to incorporate visual processing, enabling it to interpret both text and images. We use the 2024-11-20 version of the model.
  • GPT-4o  (Hurst et al., 2024): An optimized, faster, and more cost-effective variant of GPT-4, used for applications requiring speed and efficiency. We use the 2024-05-13 version of the model.
  • Gemini 2.0  (Anil et al., 2023): One of Google’s most recent models with a focus on agent capabilities supporting different input modalities such as vision, audio and text. We use Gemini-2.0-Flash.
Appendix B Puzzle Descriptions
@@
  • ATMPuzzle: The solver is presented with a bank interface which contains two options: Deposit and Withdraw. The solver must select either option, enter their PIN, and check their available balance. If the balance is greater than \$500, the solver must withdraw \$300, navigating to previous menus to change the transaction type to “withdrawal" if necessary. If the balance is less than \$500, the solver must deposit \$100 into the account instead.
@@
  • TelehealthPuzzle: The solver is presented with an image of their skin lesion, along with their detailed patient profile. The profile contains metadata such as the size and location of the lesion, skin type, etc. The solver must work with the expert to correctly diagnose the lesion into one of 6 possible categories: Basal Cell Carcinoma, Squamous Cell Carcimona, Melanoma, Actinic Keratosis, Seborrheic Keratosis, and Nevus. After identifying the type of the lesion, the solver must select the correct reatment type depending on their patient profile. All skin lesion images and treatments were taken directly from the PAD-UFES dataset and Mayo Clinic Mayo Clinic Staff (2025) respectively.
@@
  • ColorPuzzle: The solver is presented with a $\times4$ grid of colored tiles. The solver must first identify the color group with the fewest squares on a 4x4 grid and press all the squares of that color to start the module. The solver then needs to refer to a table to determine the next group to press based on the current configuration. Pressing any incorrect square results in a strike and resets the module. Non-white squares may change color after each stage. The goal is to make all squares on the grid white by following the correct sequence of groups.
@@
  • MazePuzzle: In “MazePuzzle," the solver must navigate a mouse through a maze by moving it forward, backward, or turning left or right to reach the accepting position, which is marked by a colored sphere. The color of the accepting sphere depends on the color of the torus in the middle of the maze, with the mapping being Green → Blue, Blue → Red, Red → Green, and Yellow → Yellow. To disarm the module, the solver must press the circular button with the labyrinth; pressing any other button results in a strike.
  • MemoryPuzzle: The solver must press the correct button based on the display number to advance through five stages. Incorrect presses reset the module to stage 1. Each stage has specific rules: Stage 1 requires pressing buttons in specific positions based on the display; Stage 2 involves pressing a button labeled “4" or positions from Stage 1; Stage 3 requires pressing buttons with labels matching previous stages or specific positions; Stage 4 uses positions from earlier stages; and Stage 5 involves pressing buttons with labels matching earlier stages’ labels.
  • PasswordPuzzle: The solver cycles through letters above and below each position to form a word. Each cycle displays three consecutive letters, and only one combination will match a predefined list of possible words. Once the correct word is set, the solver must press the submit button to complete the puzzle. The list of possible words includes terms like "about," "after," "great," and "write."
@@ Appendix C Additional Statistics
[…]
Table 3: Average success rate across our benchmark puzzles. For each row, the solver and expert are separate instances of the same model. “Human" model indicates a human is the solver, and the expert is a GPT-4o model. The partial success rate is calculated by averaging over 100, 30 or 10 independent runs for AI, o4-mini, or human solver respectively. The overall column is calculated by averaging across all the puzzles. We bold the human score and the best AI model in each column.
[…]
Table 4: Average mistake rate of the solver. For each row, the solver and expert are separate instances of the same model. “Human" model indicates a human is the solver, and the expert is a GPT-4o model. The partial success rate is calculated by averaging over 100, 30 or 10 independent runs for AI, o4-mini, or human solver respectively. The overall column is calculated by averaging across all the puzzles. We bold the human score and the best AI model in each column.
[…]
Table 5: Average conversation length between the solver and expert on our benchmark. For each row, the solver and expert are separate instances of the same model. “Human" model indicates a human is the solver, and the expert is a GPT-4o model. The partial success rate is calculated by averaging over 100, 30 or 10 independent runs for AI, o4-mini, or human solver respectively. The overall column is calculated by averaging across all the puzzles. We bold the human score and the best AI model in each column. We observed that the open source reasoning models often have lower conversation lengths because the solver tries actions on its own, ignoring the advice of the expert.